OpenStudy (anonymous):

∫sin^-1x dx/√ 1-x^2 = ?

hartnn (hartnn):

did you try to put \(\large x = \sin t\) ?

hartnn (hartnn):

when \(x =\sin t \\ \sin^{-1}x = t \\ dx =....?\)

OpenStudy (anonymous):

I tried to do it like this Let,1-x^2=z =>d/dx(1)-d/dx x^2=dz/dx

hartnn (hartnn):

good try, but you'll get stuck after some steps. can you please try the substitution, x= sin t now ?

OpenStudy (anonymous):

Ok thanks

hartnn (hartnn):

tell me if you get stuck with that or couldn't find the correct final answer.

OpenStudy (anonymous):

sure i will..let me try it..and i will tell u the answer..if u can check out please after that .

hartnn (hartnn):

sure :)

OpenStudy (anonymous):

let,sin^-1x=z or,d/dx sin^-1x=dz/dx or,1/√ 1-x^2=dz/dx or,dx/√ 1-x^2=dz Now,∫sin^-1x dx/√ 1-x^2=∫z dz=z^1+1/1+1+c=z^2/2+c= (sin^-1x)^2/2+c is this the answer @hartnn

hartnn (hartnn):

you're absolutely correct! good work :)

OpenStudy (anonymous):

thank you :)

hartnn (hartnn):

most welcome ^_^

OpenStudy (anonymous):

^_^